# Space spanned by characteristic exponents

### v2

@@ 3. An equidistribution theorem
over $K$. For $n\in\mathbb{Z}_{>0}$, set $p_{n}=(x_{n},\overline{x_{n}})\in{\rm Per}(F)$. Let $\Gamma$ be the Zariski closure of $\{p_{n}\mid n\in\mathbb{Z}_{>0}\}$ in $\mathbb{P}^{1}\times\mathbb{P}^{1}$, which is defined over $\overline{\mathbb{Q}}$ since the $p_{n}$ are over $\overline{\mathbb{Q}}$. As the points $p_{n}$ are pairwise distinct, we have $\dim\ (\Gamma)\geq 1$. After replacing $(p_{n})$ by a suitable subsequence, we may assume that $\Gamma$ is irreducible and that $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$, i.e., no proper subvariety of $\Gamma$ contains an infinite number of $p_{n}$. Let $\pi_{j}:\Gamma\to\mathbb{P}^{1}$ be the $j$-th projection on $\Gamma\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$ to the $j$-th factor of $\mathbb{P}^{1}\times\mathbb{P}^{1}$, for $j=1,2$. It is clear that $\pi_{1}\circ F=f\circ\pi_{1}$ and $\pi_{2}\circ F=\overline{f}\circ\pi_{2}$.
There are 2 cases to consider, i.e., $\dim(\Gamma)=2$ or $\dim(\Gamma)=1$.
When $\dim(\Gamma)=2$, then $\Gamma=\mathbb{P}^{1}\times\mathbb{P}^{1}$ and the canonical $F$-invariant probability measure on the projective complex manifold $\Gamma(\mathbb{C})=\mathbb{P}^{1}(\mathbb{C})\times\mathbb{P}^{1}(\mathbb{C})$ is $\mu:=\mu_{f}\times\mu_{\overline{f}}$, where $\mu_{f}$ and $\mu_{\overline{f}}$ are the canonical probability measures on $\mathbb{P}^{1}(\mathbb{C})$ which are $f$-invariant and $\overline{f}$-invariant, respectively. (Equivalently, $\mu_{f}$ and $\mu_{\overline{f}}$ are the unique maximal entropy measures of $f$ and $\overline{f}$, respectively.)
When $\dim(\Gamma)=1$, by the solution [8] of the dynamical Manin-Mumford problem for $F$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$, the irreducible curve $\Gamma$ is periodic under $F$. After replacing $f$ by $f^{m}$ for some suitable $m\in\mathbb{Z}_{>0}$, we may assume that the irreducible curve $\Gamma$ is $F$-invariant. Still denote by $\mu$ the canonical $F$-invariant probability measure on the projective complex analytic space $\Gamma(\mathbb{C})$. Recall that if $g:X\to X$ is a polarized surjective endomorphism on a projective variety $X$ over $\mathbb{C}$ (hence $g$ is finite), then the canonical $g$-invariant probability measure $\mu_{g}$ on the projective complex analytic space $X(\mathbb{C})$ is exactly the unique probability measure with continuous potentials on $X(\mathbb{C})$ such that $g_{*}\mu_{g}=\mu_{g}$ and $g^{*}\mu_{g}=\deg(g)^{\dim(X)}\mu_{g}$. Since $f\circ\pi_{1}=\pi_{1}\circ F$ and $\overline{f}\circ\pi_{2}=\pi_{2}\circ F$ are commuting finite morphisms, via the Tate’s limiting construction of canonical measures, it is easy to deduce that
@@
[…]
For $n\in\mathbb{Z}_{>0}$, we consider the probability measure
@@
on $\Gamma$. Then $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$. By [36, §2], there exists a unique semipositive metric over $\mathcal{L}$ making $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$ an isometry; denote $\mathcal{L}$ with this metric by $\overline{\mathcal{L}}$. The canonical measure $\mu_{\overline{\mathcal{L}}}$ associated with the metrized line bundle $\overline{\mathcal{L}}$ on $\Gamma(\mathbb{C})$ is exactly the canonical $F$-invariant probability measure $\mu$ on $\Gamma(\mathbb{C})$ considered before, since the metric of ${\overline{\mathcal{L}}}$ can be obtained by Tate’s limit process like $\mu$, see [36, §2] and [32, §3.5]. We need to check the conditions (1) (small) and (2) (generic) in Theorem 3.1. The condition (1) is trivial, since $h_{\overline{\mathcal{L}}}(\Gamma)=0$ and the height of every periodic algebraic point relative to $\overline{\mathcal{L}}$ is zero, in the setting of a polarized algebraic dynamical system (cf. [32]). For the condition (2), let $V$ be an arbitrary proper subvariety of $\Gamma$ and fix $V$. After replacing $V$ by its Galois orbit, we may assume that $V$ is defined over $K$. Then the generic condition (2) follows from Lemma 3.3 immediately. Thus the claim is true.
Let $n\in\mathbb{Z}_{>0}$, take $m\in\mathbb{Z}_{>0}$ such that $\lvert\rho_{f}(x_{n})\rvert^{m}\in K$ by the assumption, and write $l=n_{f}(x_{n})$. For every $\tau\in{\rm Gal}(\overline{\mathbb{Q}}/K)$ and $10\leq j\leq l-$ 1, we have
@@
where the first equality holds because both $\tau(p_{n})$ and $F^{j}(\tau(p_{n}))$ are in the same $F$-periodic orbit of length $l$ and the appearance $dF^{l}=d(F^{l})$ in the expressions, hence $\lvert\det(dF^{l}(F^{j}(\tau(p_{n}))))\rvert=\lvert\rho_{f}(x_{n})\rvert^{2}$. Then by the definition of $\nu_{n}$, we have
[…]
@@
[…]
Let $A\to-\infty$, by the monotone convergence theorem, we have
[…]
We will show that (4.14) leads to a contradiction, by considering the case where $\dim(\Gamma)=2$ and the case where $\dim(\Gamma)=1$ separately.
When $\dim(\Gamma)=2$, it is clear that
[…]
@@
When $\dim(\Gamma)=1$, then
[…]
@@
Therefore, $f$ must be exceptional. We have finished the proof. ∎
@@
Assume that $f$ is defined over $\overline{\mathbb{Q}}$, hence $f$ is defined over a number field $K$. Suppose that the Theorem 1.4 does not hold for $f$. Let $V$ be the $\mathbb{Q}$-span of $\chi_{f}({\rm Per}^{*}(f))$ in $\mathbb{R}$, then $\dim_{\mathbb{Q}}V<\infty$. We can take $M\in\mathbb{Z}_{>0}$ and $x_{1},\dots,x_{M}\in{\rm Per}^{*}(f)$ such that $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$ generate $V$ over $\mathbb{Q}$. By enlarging $K$, we may assume that $\lvert\rho_{f}(x_{1})\rvert,\dots,\lvert\rho_{f}(x_{M})\rvert\in K$. Then for every $z_{0}\in{\rm Per}^{*}(f)$, its characteristic exponent $\chi_{f}(z_{0})$ is a linear combination of $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$ over $\mathbb{Q}$; in other words, there exist $n\in\mathbb{Z}_{>0}$ and $n_{1},\dots,n_{M}\in\mathbb{Z}$ such that
@@
which contradicts Theorem 1.5 since $f$ is not exceptional by the assumption. ∎
@@
We will finish the whole proof of Theorem 1.4 in §6.3.
According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ in the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. Therefore, the above proof of Theorem 1.4 for $f$ defined over $\overline{\mathbb{Q}}$ includes the special case where $f$ is PCF (and non-exceptional).
@@ 5.1. Pseudo-linear algebra
Denote by ${\rm PHom}(V,W)$ the set of pseudo-morphisms from $V$ to $W$. For $f,g\in{\rm PHom}(V,W)$, we define $f+g$ to be the pair $(V_{f}\cap V_{g},f|_{V_{f}\cap V_{g}}+g|_{V_{f}\cap V_{g}})$. Then ${\rm PHom}(V,W)$ is a commutative semigroup with $+$ as the operation. We denote by 00 the pair $(V,0)$. We have $0+f=f$ for all $f\in{\rm PHom}(V,W)$. For every $a\in\mathbb{R}$, we define $af$ to be the pair $(V_{f},af)$. We note that $f+(-f)=(V_{f},0)$, which is not 00 if $V_{f}\neq V$. We have a natural embedding ${\rm Hom}(V,W)\hookrightarrow{\rm PHom}(V,W)$.
@@
A sequence $(f_{i})_{i=1}^{\infty}$ in ${\rm PHom}(V,\mathbb{R})$ is said to be an $O$-sequence if the following conditions are satisfied:
(i) $f_{i}(O)\subseteq\mathbb{R}_{\geq 0}\cup\{\infty\}$ for $i\geq 1$; (ii) for every $\lambda\in O$, the set $\{i\geq 1:f_{i}(\lambda)\neq 0\}$ is finite.
Clearly, an infinite subsequence of an $O$-sequence is still an $O$-sequence.
@@
Let $(\lambda_{i})_{i=1}^{\infty}$ be a sequence in $O$ and $(f_{i})_{i=1}^{\infty}$ be a sequence in ${\rm PHom}(V,\mathbb{R})$. We say that $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ is an upper triangle $O$-system (resp. weak upper triangle $O$-system) if the following conditions hold:
(i) $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence;
(ii) $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}$ (resp. $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}\cup\{\infty\})$ for $i\geq 1$;
(iii) $f_{j}(\lambda_{i})=0$ for $j>i\geq 1$.
Clearly, an upper triangle $O$-system is a weak upper triangle $O$-system.
@@
Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be a weak upper triangle $O$-system. Then $(\lambda_{i})_{i=1}^{\infty}$ are linearly independent over $\mathbb{R}$.
@@
Since $f_{1}(\lambda_{1})\neq 0$, we see that $\lambda_{1}\neq 0$. Then we only need to show that for all $l\geq 2$, $\lambda_{l}$ is not contained in ${\rm span}_{\mathbb{R}}\{\lambda_{i}:i\leq l-1\}$. Otherwise, $\lambda_{l}=\sum_{i=1}^{l-1}a_{i}\lambda_{i}$ for some $l\geq 2,a_{i}\in\mathbb{R},1\leq i\leq l-1$. Then $f_{l}(\lambda_{l})=\sum_{i=1}^{l-1}a_{i}f_{l}(\lambda_{i})=0$, which contradicts to our assumption. ∎
Let $\tau:V\to V$ be an involution (here we just means that $\tau^{2}=\mathrm{id}$ and $\tau$ may be the identity on $V$).
@@
Assume that $\tau(O)\subseteq O$. Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangle $O$-system. Then there exists a strictly increasing sequence $(m_{i})_{i=1}^{\infty}$ in $\mathbb{Z}_{>0}$ such that the pair $((\lambda_{m_{i}}+\tau(\lambda_{m_{i}}))_{i=1}^{\infty},(f_{m_{i}})_{i=1}^{\infty})$ is a weak upper triangle $O^{\prime}$-system, where $O^{\prime}=\{\lambda_{m_{i}}+\tau(\lambda_{m_{i}}):i\in\mathbb{Z}_{>0}\}$.
@@
Assume that we have constructed $m_{1},\dots,m_{l}$ satisfying the conditions for weak upper triangle $O^{\prime}$-systems. Since $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence and $\tau(\lambda_{1}),\cdots,\tau(\lambda_{l})\in O$, there exists $m_{l+1}>m_{l}$ such that $f_{m_{l+1}}(\tau(\lambda_{i}))=0$ for all $i=1,\dots,l$. Then for all $i=1,\dots,l$, we have
@@
We conclude the proof. ∎
By Lemma 5.4 and Lemma 5.3 we get the following result.
@@
Assume that $\tau(O)\subseteq O$. Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangle $O$-system. Then $\dim_{\mathbb{R}}{\rm span}_{\mathbb{R}}\{2^{-1}(\lambda_{i}+\tau(\lambda_{i})):i\geq 1\}=\infty$.
@@
Note that the discussion in this subsection also applies with $\mathbb{R}$ replaced by any ordered field $F$ (for example, $\mathbb{Q}$ with the usual order).

### v3

@@ 3. An equidistribution theorem
over $K$. For $n\in\mathbb{Z}_{>0}$, set $p_{n}=(x_{n},\overline{x_{n}})\in{\rm Per}(F)$. Let $\Gamma$ be the Zariski closure of $\{p_{n}\colon n\in\mathbb{Z}_{>0}\}$ in $\mathbb{P}^{1}\times\mathbb{P}^{1}$, which is defined over $\overline{\mathbb{Q}}$ since the $p_{n}$ are defined over $\overline{\mathbb{Q}}$. As the points $p_{n}$ are pairwise distinct, we have $\dim(\Gamma)\geq 1$. After passing to a subsequence of $(p_{n})$, we may assume that $\Gamma$ is irreducible and that $(p_{n})_{n=1}^{\infty}$ is generic in $\Gamma$, i.e., no proper subvariety of $\Gamma$ contains infinitely many $p_{n}$. Let $\pi_{j}:\Gamma\to\mathbb{P}^{1}$ be the $j$-th projection from $\Gamma\subseteq\mathbb{P}^{1}\times\mathbb{P}^{1}$ to the $j$-th factor, for $j=1,2$. It is clear that $\pi_{1}\circ F=f\circ\pi_{1}$ and $\pi_{2}\circ F=\overline{f}\circ\pi_{2}$.
There are 2 cases to consider, i.e., $\dim(\Gamma)=2$ and $\dim(\Gamma)=1$.
If $\dim(\Gamma)=2$, then $\Gamma=\mathbb{P}^{1}\times\mathbb{P}^{1}$, and the canonical $F$-invariant probability measure on the projective complex manifold $\Gamma(\mathbb{C})=\mathbb{P}^{1}(\mathbb{C})\times\mathbb{P}^{1}(\mathbb{C})$ is $\mu:=\mu_{f}\times\mu_{\overline{f}}$, where $\mu_{f}$ and $\mu_{\overline{f}}$ are the canonical probability measures on $\mathbb{P}^{1}(\mathbb{C})$ which are $f$-invariant and $\overline{f}$-invariant, respectively. (Equivalently, $\mu_{f}$ and $\mu_{\overline{f}}$ are the unique maximal entropy measures of $f$ and $\overline{f}$, respectively.)
If $\dim(\Gamma)=1$, by the solution [8] of the dynamical Manin-Mumford problem for $F$ on $\mathbb{P}^{1}\times\mathbb{P}^{1}$, the irreducible curve $\Gamma$ is periodic under $F$. After replacing $f$ by $f^{m}$ for a suitable $m\in\mathbb{Z}_{>0}$, we may assume that the irreducible curve $\Gamma$ is $F$-invariant. Still denote by $\mu$ the canonical $F$-invariant probability measure on the projective complex analytic space $\Gamma(\mathbb{C})$. Recall that if $g:X\to X$ is a polarized surjective endomorphism on a projective variety $X$ over $\mathbb{C}$ (hence $g$ is finite), then the canonical $g$-invariant probability measure $\mu_{g}$ on the projective complex analytic space $X(\mathbb{C})$ is exactly the unique probability measure with continuous potentials on $X(\mathbb{C})$ such that $g_{*}\mu_{g}=\mu_{g}$ and $g^{*}\mu_{g}=\deg(g)^{\dim(X)}\mu_{g}$. Since $f\circ\pi_{1}=\pi_{1}\circ F$ and $\overline{f}\circ\pi_{2}=\pi_{2}\circ F$ are commuting finite morphisms, Tate’s limiting construction of canonical measures implies that
@@
Now we deal with both cases ($\dim(\Gamma)=2$ and $\dim(\Gamma)=1$) simultaneously. We will show that $2\mathcal{L}_{f}<\int\log\lvert\det(dF)\rvert d\mu$.
For $n\in\mathbb{Z}_{>0}$, consider the probability measure
@@
on $\Gamma$. Then $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$. By [36, §2], there exists a unique semipositive metric on $\mathcal{L}$ making $(F|_{\Gamma})^{*}\mathcal{L}\cong\mathcal{L}^{\otimes d}$ an isometry; denote this metrized line bundle by $\overline{\mathcal{L}}$. The canonical measure $\mu_{\overline{\mathcal{L}}}$ associated with the metrized line bundle $\overline{\mathcal{L}}$ on $\Gamma(\mathbb{C})$ is exactly the canonical $F$-invariant probability measure $\mu$ on $\Gamma(\mathbb{C})$, since the metric of $\overline{\mathcal{L}}$} can be obtained by Tate’s limit process, similar to $\mu$; see [36, §2] and [32, §3.5]. We need to verify the conditions (1) (small) and (2) (generic) in Theorem 3.1. Condition (1) is trivial because $h_{\overline{\mathcal{L}}}(\Gamma)=0$ and the height of every periodic algebraic point relative to $\overline{\mathcal{L}}$ is zero in the setting of a polarized algebraic dynamical system (cf. [32]). For the condition (2), let $V$ be an arbitrary proper subvariety of $\Gamma$. After replacing $V$ by its Galois orbit, we may assume that $V$ is defined over $K$. Then the generic condition (2) follows immediately from Lemma 3.3. Thus the claim holds.
Let $n\in\mathbb{Z}_{>0}$, take $m\in\mathbb{Z}_{>0}$ such that $\lvert\rho_{f}(x_{n})\rvert^{m}\in K$ by the assumption, and write $l=n_{f}(x_{n})$. For every $\tau\in{\rm Gal}(\overline{\mathbb{Q}}/K)$ and $0\leq j\leq l-1$, we have
@@
where the first equality holds because both $\tau(p_{n})$ and $F^{j}(\tau(p_{n}))$ belong to the same $F$-periodic orbit of length $l$, and $dF^{l}=d(F^{l})$ in the expressions. Hence $\lvert\det(dF^{l}(F^{j}(\tau(p_{n}))))\rvert=\lvert\rho_{f}(x_{n})\rvert^{2}$. Then by the definition of $\nu_{n}$, we have
$\int\log\lvert\det(dF)\rvert d\nu_{n}=\frac{1}{l}\int\log\lvert\det(dF^{l})\rvert d\nu_{n}$
@@
$2a=2\lim\limits_{n\to\infty}\chi_{f}(x_{n})=\lim\limits_{n\to\infty}\int\log\lvert\det(dF)\rvert d\nu_{n}$
$\leq$ $\lim\limits_{n\to\infty}\int\max\{\log\lvert\det(dF)\rvert,A\}d\nu_{n}$
$=$ $\int\max\{\log\lvert\det(dF)\rvert,A\}d\mu.$
Letting $A\to$-$\infty$, by the monotone convergence theorem, we have
(4.14) $2\mathcal{L}_{f}<2a\leq\int\log\lvert\det(dF)\rvert d\mu.$
We now show that (4.14) leads to a contradiction by considering the two cases separately.
If $\dim(\Gamma)=2$, then
$\int\log\lvert\det(dF)\rvert d\mu=\int_{\mathbb{P}^{1}\times\mathbb{P}^{1}}\log\lvert\det(d(f\times\overline{f}))\rvert d(\mu_{f}\times\mu_{\overline{f}})=\mathcal{L}_{f}+\mathcal{L}_{\overline{f}}=2\mathcal{L}_{f},$
@@
If $\dim(\Gamma)=1$, then
$\int\log\lvert\det(dF)\rvert d\mu=\int_{\Gamma}\log\lvert\det(d(f\times\overline{f}))\rvert d\mu$
$=$ $\int_{\Gamma}\log\lvert\det(df)\circ\pi_{1}\rvert d\mu+\int_{\Gamma}\log\lvert\det(d\overline{f})\circ\pi_{2}\rvert d\mu$
$=$ $\int_{\Gamma}\log\lvert\det(df)\circ\pi_{1}\rvert d\frac{\pi_{1}^{*}(\mu_{f})}{\deg(\pi_{1})}+\int_{\Gamma}\log\lvert\det(d\overline{f})\circ\pi_{2}\rvert d\frac{\pi_{2}^{*}(\mu_{\overline{f}})}{\deg(\pi_{2})}$
$=$ $\int_{\mathbb{P}^{1}}\log\lvert\det(df)\rvert d\mu_{f}+\int_{\mathbb{P}^{1}}\log\lvert\det(d\overline{f})\rvert d\mu_{\overline{f}}$
@@
Therefore, $f$ must be exceptional. This completes the proof. ∎
@@
Assume that $f$ is defined over $\overline{\mathbb{Q}}$, hence $f$ is defined over some number field $K$. Suppose that Theorem 1.4 fails for $f$. Let $V$ be the $\mathbb{Q}$-span of $\chi_{f}({\rm Per}^{*}(f))$ in $\mathbb{R}$; then $\dim_{\mathbb{Q}}V<\infty$. We can take $M\in\mathbb{Z}_{>0}$ and $x_{1},\dots,x_{M}\in{\rm Per}^{*}(f)$ such that $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$ generate $V$ over $\mathbb{Q}$. By enlarging $K$, we may assume that $\lvert\rho_{f}(x_{1})\rvert,\dots,\lvert\rho_{f}(x_{M})\rvert\in K$. Then for every $z_{0}\in{\rm Per}^{*}(f)$, its characteristic exponent $\chi_{f}(z_{0})$ is a $\mathbb{Q}$-linear combination of $\chi_{f}(x_{1}),\dots,\chi_{f}(x_{M})$; in other words, there exist $n\in\mathbb{Z}_{>0}$ and $n_{1},\dots,n_{M}\in\mathbb{Z}$ such that
@@
which contradicts Theorem 1.5 since $f$ is not exceptional by assumption. ∎
@@
We will complete the whole proof of Theorem 1.4 in §6.3.
According to [6], PCF maps are defined over $\overline{\mathbb{Q}}$ within the moduli space $\mathcal{M}_{d}$ of rational maps of degree $d$, except for the family of flexible Lattès maps. Therefore, the above proof of Theorem 1.4 for $f$ defined over $\overline{\mathbb{Q}}$ includes the special case where $f$ is PCF (and non-exceptional).
@@ 5.1. Pseudo-linear algebra
Denote by ${\rm PHom}(V,W)$ the set of pseudo-morphisms from $V$ to $W$. For $f,g\in{\rm PHom}(V,W)$, we define $f+g$ to be the pair $(V_{f}\cap V_{g},f|_{V_{f}\cap V_{g}}+g|_{V_{f}\cap V_{g}})$. We denote by 0 the pair $(V,0)$. We have $0+f=f$ for all $f\in{\rm PHom}(V,W)$. Then ${\rm PHom}(V,W)$ is a commutative monoid under the operation ++ with the zero element 0. For every $a\in\mathbb{R}$, we define $af$ to be the pair $(V_{f},af)$. Note that $f+(-f)=(V_{f},0)$, which is not 0 if $V_{f}\neq V$. We have a natural embedding ${\rm Hom}(V,W)\hookrightarrow{\rm PHom}(V,W)$.
@@
A sequence $(f_{i})_{i=1}^{\infty}$ in ${\rm PHom}(V,\mathbb{R})$ is called an $O$-sequence if the following conditions hold:
  1. (i)
    $f_{i}(O)\subseteq\mathbb{R}_{\geq 0}\cup\{\infty\}$ for $i\geq 1$;
  2. (ii)
    for every $\lambda\in O$, the set $\{i\geq 1:f_{i}(\lambda)\neq 0\}$ is finite.
Clearly, any infinite subsequence of an $O$-sequence is also an $O$-sequence.
@@
Let $(\lambda_{i})_{i=1}^{\infty}$ be a sequence in $O$ and $(f_{i})_{i=1}^{\infty}$ be a sequence in ${\rm PHom}(V,\mathbb{R})$. We say that $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ is an upper triangular $O$-system (resp. weak upper triangular $O$-system) if the following conditions hold:
  1. (i)
    $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence;
  2. (ii)
    $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}$ (resp. $f_{i}(\lambda_{i})\in\mathbb{R}_{>0}\cup\{\infty\}$) for $i\geq 1$;
  3. (iii)
    $f_{j}(\lambda_{i})=0$ for $j>i\geq 1$.
Clearly, an upper triangular $O$-system is a weak upper triangular $O$-system.
@@
Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be a weak upper triangular $O$-system. Then $(\lambda_{i})_{i=1}^{\infty}$ are linearly independent over $\mathbb{R}$.
@@
Since $f_{1}(\lambda_{1})\neq 0$, we have $\lambda_{1}\neq 0$. It remains to show that for all $l\geq 2$, $\lambda_{l}$ is not contained in ${\rm span}_{\mathbb{R}}\{\lambda_{i}:i\leq l-1\}$. Suppose otherwise that $\lambda_{l}=\sum_{i=1}^{l-1}a_{i}\lambda_{i}$ for some $l\geq 2$ and $a_{i}\in\mathbb{R}$, $1\leq i\leq l-1$. Then $f_{l}(\lambda_{l})=\sum_{i=1}^{l-1}a_{i}f_{l}(\lambda_{i})=0$, contradicting our assumption. ∎
Let $\tau:V\to V$ be an involution (here we only assume that that $\tau^{2}=\mathrm{Id}$; $\tau$ may be the identity on $V$).
@@
Assume that $\tau(O)\subseteq O$. Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangular $O$-system. Then there exists a strictly increasing sequence $(m_{i})_{i=1}^{\infty}$ in $\mathbb{Z}_{>0}$ such that the pair $((\lambda_{m_{i}}+\tau(\lambda_{m_{i}}))_{i=1}^{\infty},(f_{m_{i}})_{i=1}^{\infty})$ is a weak upper triangular $O^{\prime}$-system, where $O^{\prime}=\{\lambda_{m_{i}}+\tau(\lambda_{m_{i}}):i\in\mathbb{Z}_{>0}\}$.
@@
Assume that we have constructed $m_{1},\dots,m_{l}$ satisfying the conditions for weak upper triangular $O^{\prime}$-systems. Since $(f_{i})_{i=1}^{\infty}$ is an $O$-sequence and $\tau(\lambda_{1}),\dots,\tau(\lambda_{l})\in O$, there exists $m_{l+1}>m_{l}$ such that $f_{m_{l+1}}(\tau(\lambda_{m_{i}}))=0$ for all $1\leq i\leq l$. Then for all $i=1,\dots,l$, we have
@@
This completes the proof. ∎
Combining Lemma 5.4 and Lemma 5.3, we obtain:
@@
Assume that $\tau(O)\subseteq O$. Let $((\lambda_{i})_{i=1}^{\infty},(f_{i})_{i=1}^{\infty})$ be an upper triangular $O$-system. Then $\dim_{\mathbb{R}}{\rm span}_{\mathbb{R}}\{2^{-1}(\lambda_{i}+\tau(\lambda_{i})):i\geq 1\}=\infty$.
@@
The discussion in this subsection also applies when $\mathbb{R}$ is replaced by any ordered field $F$ (for example, $\mathbb{Q}$ with the usual order).